PLEASE HELP ME!
Apply distributive property, simplify combining like terms.
2 ( 3a - 5 ) + 2a = ??
My answer that i got was a x 2 but i think its wrong PLEASE HELP!

Answer by venugopalramana(3286)

2*( 3a - 5 ) + 2a = ??...distributive property here is that multiplication distributes over subtraction..that is
2*3a-2*5+2a=6a-10+2a=6a+2a-10=8a-10